OpenStudy (anonymous):

A country's population in 1993 was 94 million. In 1999 it was 99 million. Estimate the population growth formula. Round your answer to the nearest million. P=Ae^kt

OpenStudy (phi):

I would let t=0 (the "start time" ) correspond to 1993 population at the start i.e. at t=0 (1993) is 94 At t=0, we can write \[ A e^{k\cdot 0} = 94 \\ A e^0 = 94\] remember anything to the zero power is 1. that is e^0 =1, so \[ A e^0 = 94 \\ A= 94\] in 1999, which is 6 years past 1993, t=6 we know at t=6 (i.e. 1999), P = 99 write \[ 94 e^{6k} = 99\]

OpenStudy (phi):

we now can solve for k. first divide both sides by 94 \[ 94 e^{6k} = 99 \\ e^{6k} = \frac{99}{94} \] "take the natural log" of both sides \[ \ln\left(e^{6k} \right) = \ln\left( \frac{99}{94} \right) \\ 6k = \ln\left( \frac{99}{94} \right) \\ k = \frac{1}{6}\ \ln\left( \frac{99}{94} \right) \] we will need to use a calculator to find a decimal value for k

OpenStudy (phi):

we will find k= 0.00863751131 \[ P= 94 e^{0.00864 t} \]